John's granddaughter's age is about as many days as his daughter's age is in weeks, and his granddaughter's age is as many months as he is in years. The ages of John, his daughter, and his granddaughter, together are 120 years. How old is each person in years?
Answer by macston(5194)

G=grand-daughter's age; D=daughter's age=7G; J=John's age=12G
G+D+J=120yrs Substitute for D and J, solve for G
G+7G+12G=120yrs
20G=120yrs
G=6yrs ANSWER 1: The grand-daughter is 6 years old.
D=7G=7(6yrs)=42yrs ANSWER 2: The daughter is 42 years old.
J=12G=12(6yrs)=72yrs ANSWER 3: John is 72 years old.
CHECK
J+D+G=120yrs
72yrs+42yrs+6yrs=120yrs
120yrs=120yrs